OpenStudy (anonymous):

Solve, using the substitution method. j + k = 3 j – k = 7 A. The solution is (8, 1) B. The solution is (5, –2) C. There is no solution. D. There are an infinite number of solutions.

OpenStudy (anonymous):

B is the correct option.... j+k=3............(1) j-k=7 ............(2) Add (1) n (2), you get: 2j=10 => j = 5; put back j in (1) or (2) k=-2